# A Closer Look into Mixture-of-Experts in Large Language Models

### v1

@@ Abstract
Mixture-of-experts (MoE) is gaining increasing attention due to its unique properties and remarkable performance, especially for language tasks. By sparsely activating a subset of parameters for each token, MoE architecture could increase the model size without sacrificing computational efficiency, achieving a better trade-off between performance and training costs. However, the underlying mechanism of MoE still lacks further exploration, and its modularization degree remains questionable. In this paper, we make an initial attempt to understand the inner workings of MoE-based large language models. Concretely, we comprehensively study the parametric and behavioral features of three recent MoE-based models and reveal some intriguing observations, including (1) Neurons act like fine-grained experts. (2) The router of MoE usually selects experts with larger output norms. (3) The expert diversity increases as the layer increases, while the last layer is an outlier. Based on the observations, we also provide suggestions for a broad spectrum of MoE practitioners, such as router design and expert allocation. We hope this work could shed light on future research on the MoE framework and other modular architectures. Code is available at https://github.com/kamanphoebe/Look-into-MoEs.
@@ 1 Introduction
The advent of Large Language Models (LLMs) revolutionized the field of Natural Language Processing. LLMs researchers are continually pushing the boundaries of Language Models by scaling up both model size and the column of training data, significantly enhancing the capabilities of these models. This escalation in training cost and complexity necessitates innovative solutions to better balance between pre-training efficiency and model performance. One emerging solution to this end is the Mixture-of-Experts (MoE) Shazeer et al. (2017) architecture. The MoE framework facilitates the computational efficiency of the model by dynamically routing inputs to a subset of experts, allowing for substantial model scaling while maintaining training costs and leading to numerous influential advancements in the field Reid et al. (2024); Jiang et al. (2024); Dai et al. (2024); Team (2024).
@@
Beyond efficiency, another attractive trait of MoE architecture is its modular design and learning paradigm. This modularization could enable more flexible and potentially more generalizable handling of diverse data and tasks within a single MoE model by assigning them to more specialized experts. Despite its widespread adoption, it remains an open question whether existing MoE-based LLMs truly leverage this modularity in their knowledge distribution or model behaviors across different experts. In other words, is MoE a simple ensemble of homogeneous experts or a modular combination of heterogeneous experts? Answering the above question comprehensively is non-trivial. Therefore, in this paper, we take the first step by investigating three recent MoE-based LLMs (Mixtral 8x7B Jiang et al. (2024), DeepSeekMoE Dai et al. (2024), and Grok-$1$https://github.com/xai-org/grok-1) from two critical perspectives: model parameters and model behaviors. We aim to explore common and distinct features and behaviors among different experts, further shedding light on the inner mechanisms of MoE-based models.
@@
Specifically, we examine the correlation between experts’ parameters, gates, and their output features given text inputs. Before diving into deeper analyses, we briefly summarize some of our empirical conclusions (detailed in Sec 6) and observations:
@@
  • Neurons in the Feed-Forward Network (FFN) layer are fine-grained experts. Both the gate embedding and the gate projection matrix for experts perform the choosing operation: the former determines the expert selection while the latter controls the neuron activation. Meanwhile, we observe that their similar heat maps exhibit correlation, revealing that from the gate projection perspective, the expert neurons can probably be regarded as “tiny” experts with one neuron.
@@
  • Increasing the number of experts in deeper layers but reducing it in the last layer. We observe that the similarities between experts’ parameters and outputs continuously decrease with the increase in layer number while suddenly increasing in the last layer.
@@
  • Employing the norm as the routing mechanism is a reasonable choice. For Mixtral 8x7B and DeepSeekMoE, we find that the gate usually selects experts with larger output norms.
@@
  • When analyzing the correlation between experts, measuring the similarities between weight matrices is, to some extent, equivalent to measuring the averaged similarities of expert outputs.
@@
  • Compared with some special initialization schemes, training MoE from scratch is more likely to facilitate expert diversity. This stems from the observations that stronger correlations (e.g., higher similarities) exist between Mixtral experts’ parameters and behaviors, while DeepSeekMoE and Grok-1, which are trained from scratch, do not show such correlations.
@@ 3 Overview
Our experiments are conducted on several open-source MoE models, namely Mixtral 8x7B, DeepSeekMoE, and Grok-1, which demonstrate impressive performance across various domains. To further study the similarities and differences between a standard transformer and a MoE model, we include Mistral 7B Jiang et al. (2023) as one of our investigated models. Some basic information about these models and the abbreviations of the model names used throughout our paper are summarized in Tab 1. The analysis is divided into two sections, focusing on the model parameters (static) and the model behaviors with text inputted (dynamic). Unless otherwise stated, cosine similarity is employed for all experiments involving similarity measurement.
@@ 4 Analysis of Static Parameters
From a high-level perspective, the knowledge a model learns is encoded in its parameters. Hence, investigating weight matrices is a nartural way to examine a model. In this section, we study the correlation between parameters of: (i) MoE experts (and FFNs for Mistral), (ii) gate embeddings, which are two vital components of MoE.
@@ 4.1 Weight Matrices of Experts
MoE models replace FFNs in standard transformers with experts. Thus, following Geva et al. (2020); Qiu et al. (2024), the projection matrices of the experts can be regarded as keys and values: the column vectors of $W_{\text{down}}$ represent the possible outputs; the row vectors of $W_{\text{up}}$ produce the weights for each possible output; the row vectors of $W_{\text{gate}}$ decide whether to activate the corresponding neurons. Therefore, investigating the weight matrices is a straightforward way to understand the experts’ behaviors. To study from different views, we analyze both the matrix and the neuron levels.
@@ 4.1.1 Matrix-level
In this part, we explore the similarity of weight matrices between all experts in every layer for the three projection matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$. The similarity is calculated based on the flattened matrices vectors and illustrated in Fig. 1. We denote “F” as the Mistral FFN and “SE” as the DeepSeek shared expert in all the figures. Note that the figures of different models do not share the same color-bar.
Common [^2] The observations shared by all of our investigated models are written in the Common part.. The heat maps of the three matrices show similar patterns. Since the weight matrices are large, directly flattening them gives high-dimension vectors. We employ principal components analysis (PCA) to convert the flattened vectors to two-dimensional space. The resulting figures show that the expert distribution across the three weight matrices is also generally alike for Mixtral and DeepSeek. More details about the PCA results are presented in Append A.1.
Mixtral and Mistral. The cosine similarities between Mixtral experts ($S_{\text{ee}}$) mostly lie between 0.2 to 0.4, while the similarities between the experts and the Mistral FFN ($S_{\text{ef}}$) are about 0.6. Yet the values tend to be lower for deep (22th-30th) layers. A “dark cross” can be observed in some layers and corresponds with the outliers in the 2D space projected by PCA, indicating that the corresponding expert is relatively distinct from the others. Interestingly, this cross appears in Expert 3 the most frequently, which suggests that this expert might have learned some special attributes. It is worth noting that the cross usually passes through the entire heat map, including the last row of FFN. Hence, when a Mixtral expert differs from other experts, it is also less similar to the FFN of Mistral.
DeepSeek and Grok. Since the shared experts of DeepSeek are implemented as a single MLP block with hidden size larger than the routed experts, their flattened vectors cannot be directly compared and thus we omit the shared experts for this experiment. Fig. 1 demonstrates that the similarities between the DeepSeek routed and Grok experts are close to zero. While Mixtral’s training method remains unrevealed, DeepSeek and Grok are known to be trained from scratch. This experiment shows that Mixtral might be trained by some special schemes, leading to less diverse experts training from scratch Wu et al. (2022).
@@
Figure 1: Matrix-level similarity heat maps of expert weight matrices. Each layer contains three heat maps, corresponding to $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, respectively. The tick numbers refer to expert indices.
@@ 4.1.2 Neuron-level
In Sec 4.1, we measure the parameter similarity between experts in matrix level. However, the calculation of cosine similarity is position-related. Assume that the neurons of two experts are similar but in different order, the similarity of their weight matrices will be much lower than expected. To address this, we propose two approaches to investigate the correlation in neuron level: averaging and reordering. Averaging simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices and then calculates the similarity of the resulting vectors across experts. For reordering, we apply the Jonker-Volgenant algorithm Jonker and Volgenant (1988), typically used for solving linear assignment problems, to find the optimal order of neurons to maximize the cosine similarity of the two experts. We describe the results of the reordering method below and leave the averaging details in Append B. Additionally, the projection of neurons in low-dimensional spaces using PCA can be found in Append A.2. Due to the heavy computation, we only pick several layers for the reordering calculation. Note that the matrices are reordered separately. We measure the Kendall’s $\tau$ coefficient between the index sequences before and after reordering, whose value increases when two sequences exist strong agreement. Tab. 2 depicts the common similarity growth after reordering and the averaged Kendall’s coefficient $\bar{\tau}$ over the selected layers. The order of Mixtral neurons hardly changes (large $\tau$) and hence almost unchanged similarities. Despite the vast similarity growth for DeepSeek and Grok after the reordering, their overall values are still about 1e-2.
@@ 4.2 Gate Embedding
The gate in the chosen MoE models is implemented as a linear layer with embedding size $\mathbb{R}^{n_{\text{exp}}}\times\mathbb{R}^{d_{\text{hid}}}$, where $n_{\text{exp}}$ represents the number of experts. The gate is another main component of MoE besides the experts, and thus, it is crucial to study its attributes to better understand how MoE works. In addition, since each row vector in its embedding matrix determines the corresponding expert selection, some correspondence may exist between the embedding matrix and the expert weights.
To investigate this, we measure the similarities between the gate embedding vectors. For computation simplicity, we compare them with the averaged heat maps (instead of the reordering results) in Append B, and the qualitative analyses are detailed in Append C. Specifically, we found that, for all three MoE models, the patterns in the heat maps of gate vectors and expert neurons in $W_{\text{gate}}$ are partially alike in some layers (i.e., the same coordinates in the two heat maps exhibit relatively higher/lower values simultaneously).
Therefore, we further conduct a quantitative analysis of their similarity values. In particular, we perform linear regression on the paired similarity dataset $(X,Y)$, where $X$ and $Y$ denote the similarities of the gate vectors and one of the weight matrices $W_{\text{up}}$, $W_{\text{gate}}$, $W_{\text{down}}$, respectively. Tab 3 describes the square of Pearson correlation coefficients averaged over all layers ($R_{\text{avg}}^{2}$) and Tab 4 in the Appendix lists the Pearson correlation coefficient ($R$) of every layer. As shown in Tab 3, the correlation between the similarities of the gate vectors and of $W_{\text{gate}}$ is significantly stronger than $W_{\text{up}}$ and $W_{\text{dwon}}$. For the $(X,Y_{\text{gate}})$ pair, although Mixtral and DeepSeek have similar $R_{\text{avg}}^{2}$, $R^{2}$ of Mixtral fluctuates between 0.1 and 0.7 while $R^{2}$ of DeepSeek stays near to 0.4. Furthermore, we can see from Tab 4 that $(X,Y_{\text{gate}})$ of both Mixtral and DeepSeek show positive correlations, whereas $(X,Y_{\text{gate}})$ of Grok turn to negative correlations starting from intermediate (>25th) layers. We note that the function of the gate embedding and $W_{\text{gate}}$ is alike: the former determines the expert selection while the latter is responsible for choosing neurons to activate. Therefore, they may learn similar knowledge to perform the choosing operation reasonably, hence the observed correlation.
@@
Table 3: Square of Pearson correlation coefficients averaged over all layers ($R_{\text{avg}}^{2}$) for three paired dataset.
@@ 4.3 Summary
Here, we conclude the inspiring observations in the analysis of static parameters: (1) Mixtral might contain expert(s) with special attributes. Dark crosses can be frequently found in Fig. 1. (2) The similarities of DeepSeek and Grok expert weight matrices are generally lower than those of Mixtral. As mentioned in Sec 4.1.1, the matrix-level similarities of DeepSeek and Grok experts are usually zero, whereas the Mixtral expert similarities reach about 0.3 on average. (3) Different experts’ weights are less similar in deep layers. This can be observed from the Mixtral heat maps in Fig. 1. (4) $W_{\text{up}}$, $W_{\text{down}}$, and $W_{\text{gate}}$, share similar patterns in their similarity heat maps (Fig. 1). (5) The similarities of the gate embeddings and of $W_{\text{gate}}$ show either positive or negative association. Tab 3 depicts the $R_{\text{avg}}^{2}$ values, where the gate embeddings and $W_{\text{gate}}$ pair achieves the highest for all three models.
@@ 5 Analysis of Dynamic Behaviours
Figure 2: Similarity heat maps of expert output features using the short sequence. The top $k$ experts for each token are shown on top of each heat map. The tick numbers refer to expert indices.
@@
Figure 3: Averaged similarity heat maps of expert output features using the long sequence, plotted along with the matrix-level similarity heat maps. The tick numbers refer to expert indices.
The previous experiments examine the MoE models via their parameters, which do not involve any input. In this section, we feed text sequences into the MoE models to further study their actual behaviours given various inputs. Specifically, we analyze the outputs of the experts and gates.

### v2

@@ Abstract
Mixture-of-experts (MoE) is gaining increasing attention due to its unique properties and remarkable performance, especially for language tasks. By sparsely activating a subset of parameters for each token, MoE architecture could increase the model size without sacrificing computational efficiency, achieving a better trade-off between performance and training costs. However, the underlying mechanism of MoE still lacks further exploration, and its modularization degree remains questionable. In this paper, we make an initial attempt to understand the inner workings of MoE-based large language models. Concretely, we comprehensively study the parametric and behavioral features of three popular MoE-based models and reveal some intriguing observations, including 1) Neurons act like fine-grained experts; 2) The router of MoE usually selects experts with larger output norms; 3) The expert diversity increases as the layer increases, while the last layer is an outlier, which is further validated by an initial experiment. Based on the observations, we also provide suggestions for a broad spectrum of MoE practitioners, such as router design and expert allocation. We hope this work could shed light on future research on the MoE framework and other modular architectures. Code is available at https://github.com/kamanphoebe/Look-into-MoEs.
@@ 1 Introduction
The advent of Large Language Models (LLMs) revolutionized the field of Natural Language Processing. LLM researchers are continually pushing the boundaries of Language Models by scaling up both model size and the column of training data, significantly enhancing the capabilities of these models. This escalation in training cost and complexity necessitates innovative solutions to better balance between pre-training efficiency and model performance. One emerging solution to this end is the Mixture-of-Experts (MoE) Shazeer et al. (2017) architecture. The MoE framework facilitates the computational efficiency of the model by dynamically routing inputs to a subset of experts, allowing for substantial model scaling while maintaining training costs and leading to numerous influential advancements in the field Reid et al. (2024); Jiang et al. (2024); Dai et al. (2024); Team (2024).
@@
Beyond efficiency, another attractive trait of MoE architecture is its modular design and learning paradigm. This modularization allows for flexible and potentially more generalizable handling of diverse data and tasks within a single model by assigning them to specialized experts. Despite its widespread adoption, it remains an open question whether current MoE-based LLMs truly leverage this modularity in knowledge distribution and expert behaviors. In other words, is MoE a simple ensemble of homogeneous experts or a modular combination of heterogeneous experts? Answering this question comprehensively is non-trivial. Therefore, in this paper, we take the first step by investigating three popular MoE-based LLMs (Mixtral 8x7B Jiang et al. (2024), DeepSeekMoE Dai et al. (2024), and Grok-$1$https://github.com/xai-org/grok-1) from two critical perspectives: model parameters and model behaviors. We aim to explore common and distinct features and behaviors among different experts, further shedding light on the inner mechanisms of MoE-based models.
@@
Specifically, we examine the correlation between experts’ parameters, gates, and their output features given text inputs. Before diving into deeper analyses, we briefly summarize some of our empirical conclusions (detailed in § 6) and observations:
@@
  • Neurons in the Feed-Forward Network (FFN) layer are fine-grained experts. Both the gate embedding and the gate projection matrix of experts perform the choosing operation: the former determines the expert selection while the latter controls the neuron activation. We observe that the similarity heat maps exhibit correlations, suggesting that, from the perspective of the gate projection, the expert neurons can be considered as “tiny” experts, each represented by a single neuron.
@@
  • Increasing the number of experts in deeper layers while reducing it in the last layer. This is experimented in Fig. 5. Our observations indicate that the similarities between the parameters and outputs of the experts consistently decrease with increasing layer number, followed by a sudden increase in the last layer.
@@
  • Using the norm as the routing mechanism is a reasonable choice. For both Mixtral 8x7B and DeepSeekMoE, we observe that the gate typically selects experts with larger output norms.
@@
  • When analyzing the correlation between experts, measuring the similarities between weight matrices is, to some extent, equivalent to assessing the average similarities of expert outputs.
@@
  • Training MoE from scratch promotes greater expert diversity than specific initialization schemes. This stems from the observations that stronger correlations (e.g., higher similarities) between parameters and behaviors in Mixtral experts. In contrast, DeepSeekMoE and Grok-1, which are trained from scratch, do not show these correlations.
@@ 3 Overview
Our experiments are conducted on several open-source MoE models, namely Mixtral 8x7B, DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, they exhibit complementary characteristics across several key attributes, enabling a robust comparative analysis using control variables. Details are discussed in Append A. To further study the similarities and differences between a standard transformer and a MoE model, we include Mistral 7B Jiang et al. (2023) as one of our investigated models. Basic information about these models, along with the abbreviations used throughout our paper, is summarized in Tab. 1 and Tab. 4. The analysis is divided into two sections: one focusing on model parameters (static) and the other on model behaviors in response to text input (dynamic). Unless otherwise stated, cosine similarity is employed for all experiments involving similarity measurements.
@@ 4 Analysis of Static Parameters
From a high-level perspective, a model’s knowledge is encoded in its parameters, making the investigation of weight matrices a natural approach. In this section, we study the correlation between the parameters of: i) MoE experts (and FFNs for Mistral), and ii) gate embeddings; which are two vital components of the MoE architecture.
@@ 4.1 Weight Matrices of Experts
MoE models replace FFNs in standard transformers with experts. Following Geva et al. (2020); Qiu et al. (2024b), the projection matrices of the experts can be regarded as keys and values: the column vectors of $W_{\text{down}}$ represent potential outputs; the row vectors of $W_{\text{up}}$ produce weights for each possible output; the row vectors of $W_{\text{gate}}$ determine whether to activate the corresponding neurons. Thus, examining the weight matrices provides a straightforward way to understand the expert behaviors. We analyze both the matrix and neuron levels to gain insights from different perspectives.
@@ 4.1.1 Matrix-level
In this part, we explore the similarity of weight matrices among all experts in each layer for the three projection matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$. The similarity is calculated based on the flattened matrices and is illustrated in Fig. 1. We denote “F” as the Mistral FFN and “SE” as the DeepSeek shared expert. Note that the figures for different models do not share the same color bar.
Common  [^2] The observations shared by all of our investigated models are written in the Common part.. The heat maps of the three matrices exhibit similar patterns. Directly flattening the large weight matrices leads to high-dimensional vectors, so we use principal components analysis (PCA) to reduce these vectors to two-dimensional space. The resulting figures also show that, for Mixtral and DeepSeek, the expert distribution across the three weight matrices is generally comparable. Details on the PCA results are presented in Append C.1.
[…]
DeepSeek and Grok. The shared experts of DeepSeek are implemented as a single MLP block with a larger hidden size than the routed experts, preventing direct comparison of their flattened vectors; thus, we omit them from this experiment. Fig. 1 demonstrates that the similarities between the DeepSeek routed experts and Grok experts are close to zero. While Mixtral’s training method remains unrevealed, it is known that DeepSeek and Grok are trained from scratch. This suggests that Mixtral may have been trained using special schemes, resulting in less diverse experts compared to those trained from scratch Wu et al. (2022).
@@
Figure 1: Matrix-level similarity heat maps of expert weight matrices. Each layer contains three heat maps, corresponding to $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, respectively. The tick numbers refer to the expert indices.
@@ 4.1.2 Neuron-level
In § 4.1, we measure the parameter similarity between experts at the matrix level. However, the calculation of cosine similarity is position-dependent. If the neurons of two experts are similar but in different orders, the similarity of their weight matrices will be significantly lower than expected. To address this, we propose two approaches to investigate the correlation at the neuron level: averaging and reordering. Averaging simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices, and then calculates the similarity of the resulting vectors across experts. For reordering, we apply the Jonker-Volgenant algorithm Jonker and Volgenant (1988), which is typically used for solving linear assignment problems, to find the optimal order of neurons that maximizes the cosine similarity between the two experts. We describe the results of the reordering method below and provide the details of the averaging approach in Append D. Additionally, the projection of neurons into low-dimensional spaces using PCA can be found in Append C.2. Due to the heavy computation, we only select several layers for the reordering calculation. Note that the matrices are reordered separately. We measure Kendall’s $\tau$ coefficient between the index sequences before and after reordering, whose value increases when the two sequences exist strong agreement. Tab. 2 depicts the common similarity growth after reordering and the average Kendall’s coefficient $\bar{\tau}$ over the selected layers. The order of Mixtral neurons changes little (resulting in a large $\tau$), and hence nearly unchanged similarities. Despite the substantial similarity increase for DeepSeek and Grok after reordering, their overall values remain around 1e-2.
@@ 4.2 Gate Embedding
The gate in the chosen MoE models is implemented as a linear layer with an embedding size of $\mathbb{R}^{n_{\text{exp}}}\times\mathbb{R}^{d_{\text{hid}}}$, where $n_{\text{exp}}$ is the number of experts. The gate serves as a crucial component of MoE, making it essential to study its attributes to understand MoE functionality better. In addition, since each row vector in the embedding matrix determines expert selection, some correspondence may exist between the embedding matrix and the expert weights.
To investigate this, we measure the similarities between the gate embedding vectors. For computational simplicity, we compare them with the average heat maps (instead of the reordering results) presented in Append D, with qualitative analyses detailed in Append E. Specifically, we found that, for all three MoE models, the patterns in the heat maps of gate vectors and expert neurons in $W_{\text{gate}}$ are partially alike in some layers (i.e., the same coordinates in both heat maps exhibit relatively higher or lower values simultaneously).
[…]
@@
Table 3: Average square of Pearson correlation coefficients over all layers ($R_{\text{avg}}^{2}$) for three paired dataset.
@@ 4.3 Summary
Here, we conclude the key observations from the analysis of static parameters: i) Mixtral might contain expert(s) with unique attributes, as evidenced by the frequent presence of dark crosses in Fig. 1. ii) The similarities of DeepSeek and Grok expert weight matrices are generally lower than those in Mixtral. As mentioned in § 4.1.1, the matrix-level similarities of DeepSeek and Grok experts are typically close to zero, whereas Mixtral expert similarities average around 0.3. iii) The weights of different experts become less similar in deeper layers, as observed in the Mixtral heat maps in Fig. 1. iv) $W_{\text{up}}$, $W_{\text{down}}$, and $W_{\text{gate}}$, share similar patterns in their similarity heat maps (Fig. 1). v) The similarities of the gate embeddings and of $W_{\text{gate}}$ show either positive or negative association. Tab. 3 depicts the $R_{\text{avg}}^{2}$ values, where the pairing of gate embeddings and $W_{\text{gate}}$ achieves the highest correlation across all three models.
@@ 5 Analysis of Dynamic Behaviours
Figure 2: Similarity heat maps of expert output features using the short input. The top $k$ experts for each token are shown on top of each heat map. The tick numbers refer to the expert indices.
@@
Figure 3: Average similarity heat maps of expert output features using the long input, plotted along with the matrix-level similarity heat maps. The tick numbers refer to the expert indices.
The previous experiments examine the MoE models via their parameters, without involving any input. In this section, we feed text sequences into the MoE models to further study their actual behaviours given various inputs. Specifically, we analyze the outputs of the experts and gates.
@@
We utilize a short input and a long input for the experiments in this section. For the short input, we employ the first few words of the input from another MoE-related work Cai et al. (2024)  [^3] The specific tokens are ¡s¿, As, an, open, source, alternative, to, where the start of the sentence symbol ¡s¿ does not applicable for the Grok tokenizer.. For the long input, we adopt 10 sequences from the test set of the WikiText-103 Merity et al. (2016) dataset, totaling approximately 1100 tokens. The sequences in WikiText-103 cover a variety of domains, with the 10 sequences we used spanning topics such as music, design, and construction. To ensure the robustness of our findings, we repeat experiments requiring the long input (§ 5.1, § 5.2) using additional datasets with over 80K tokens. Observations align with the results described in the main context, even with subject-specific datasets like math and code. See Append F for dataset details.
We also conduct experiments for analyzing intermediate states of experts and routing patterns. Due to the page limit, these experiments are presented in Append H and Append I, respectively.